### v1

# Effective Note-taking and its Impact on Learning Undergraduate Introductory Physics Courses

Abstract
Taking notes during lectures is one of the required skills, among many others, that students need (i) to master the topic covered in the lecture, (ii) to actively engage in the learning process with minimal to no distractions, (iii) to retain learned knowledge and skills for a longer time, and (iv) in securing better grade. To learn the role of note-taking in learning undergraduate-level introductory algebra-based physics courses, we present a comparative study of students’ achievement in two minority-serving universities, namely, Fayetteville State University and The University of North Carolina at Pembroke (i) when students were taught effective note-taking strategies, motivated them to prepare notes and let them use their self-prepared notes in timed assignments and exams versus (ii) no note-taking scheme was implemented, keeping all other conditions the same. The no-note-taking scheme was used in the first introductory algebra-based physics course, while the note-taking scheme was used in the second introductory algebra-based physics course. Students’ scores in the final exam and aggregated grade accounting for all grading components are taken as measurement tools. Analysis shows that effective note-taking helps students secure better letter grades.
I Introduction
Based on the three-stage model of memory for information flow [1], namely, (i) sensory registers via visual, auditory, haptic, and so on (ii) short-term storage of information from temporary working memory and (iii) long-term storage of information, learning can be perceived as encoding information permanently in the long-term memory [2]. Sensory organs first perceive any information obtained from their source and then pass it to sensory memory. The information in sensory memory is transferred to working memory, also called short-term memory, and remains only temporarily or lost in no time. However, various control processes such as sufficient rehearsal, coding, decision-making, information retrieval, and retrieval strategies help to encode information in long-term storage and retain it there, which is otherwise lost or forgotten [1]. Learning new concepts, generating new ideas, acquiring skills, then storing them in long-term memory, and, when needed, retrieving them effortlessly and applying them is the primary goal of learning.
An effective learning process requires good organization of learning materials and attentiveness while learning. Students cannot sustain long-term attention to learning materials if activities that allow them to actively engage and contribute are omitted. In a typical lecture with no activities, students’ attentiveness peaks ( 70%) at around 10 minutes and declines after that, reaching a minimum ( 20%) at around 20 minutes [3, 4]. To maintain students’ attentiveness at the maximum possible level throughout the lecture, an instructor can incorporate various activities, visuals, simulations, and other engaging elements at regular intervals. Active processing helps move information from working memory to long-term memory [5], and taking notes is one way to do this. In this study, we explore how learning can be improved by encouraging effective note-taking and adding activities during lectures.
Effects of note-taking on learning a secondary language have been extensively studied, with the literature showing that it helps in learning the second language (see a recent review in Ref. [6]). In learning the second language vocabulary, Jin and Webb reported that vocabulary learning gains in the range of 12.5% to 23.0% [7]. Authors in Ref. [8] presented that note-taking has even better impacts on second language listening comprehension for high school students. Campana conducted a study  [9] on the effectiveness of using guided notes in high school physics by dividing the class into two groups. Therein, the author observed that the sections that were completed and used guided notes performed better academically.
To the best of our knowledge, no study has examined the role of effective note-taking in learning algebra-based undergraduate physics courses, as we presented in this research work to compare our findings; however, some works in the related science, technology, engineering, and mathematics (STEM) fields are available. Anastasiou et al. studied the effectiveness of concept maps in science learning [10]. Therein the authors used data from 55 studies published between 1980 and 2020, including peer-reviewed journals and dissertations which covered various teaching settings, conditions, and methods and divulged that the concept maps were associated with increased science learning, but there was a significant heterogeneity in most subsets of data. In one of the studies made by Shi et al on 42 college students’ learning achievement and cognitive load in a 6-week lecture-based computer network course, the authors found that note taking helped in learning   [11]. In this regard, the authors demonstrated that among the three different note-taking styles observed, namely collaborative note-taking, laptop note-taking, and traditional longhand note-taking, collaborative note-taking was found to improve learning achievement. Recently, Fanguy studied the effects of collaboratively generated notes by students versus instructor-provided notes on recalling and academic writing performance and found that student generated collaborative notes by students were more effective inenhancing academic writing, as well as in applying and developing related skills [12].
With the current technological advancement in the electronic industry, students are likely to suffer from many digital distractions in the learning environment, which interfere badly with learning [13]. In search of a better solution to minimize distraction in learning and increase attentiveness, we here present a case study on the role of effective note-taking in learning undergraduate physics courses in two minority-serving institutions Fayetteville State University (FSU) and The University of North Carolina at Pembroke (UNCP) located in the Sandhills region of North Carolina.
II Effective Note-taking
Note-taking significantly enhances deeper understanding, leading to metacognition, as writing notes helps the brain actively process and encode information, link ideas, integrate them with new knowledge, strengthen memory pathways, provide opportunities for reviewing, self-testing, and encourage reflection [14]. Note-taking offers students an elaborative rehearsal, supports analytical and critical thinking, and facilitates repetition and paraphrasing of the information, thereby enhancing their long-term memory [15]. Training students in note-taking skills, encouraging them to explore further, and helping them become effective notetakers are crucial. Taking an example of an undergraduate algebra-based physics course, the survey was conducted based on the hypothesis that note-taking helps in the learning process.
Note-taking is a crucial aspect of effective learning as it forces learners to engage with the learning material actively. It helps reinforce understanding and retention, encode information into long-term memory, provide opportunities for repetition and rehearsal, and organize information in a structured manner, making it easier to follow the logical flow of concepts. The preferred way of learning or approaching a task varies from person to person based on how the mind receives and processes information from its senses. Learners’ productivity, creativity, problem-solving skills, decision-making capability, quality of learning, and achievement can be increased and speed up by properly recognizing and understanding the learners’ preferred learning style [16]. Notes can be personalized to fit an individual’s learning styles, integrating mnemonics, incorporating annotations and highlights that give the full picture, making a good sense of the concepts to the note-taker [17]. Well-prepared notes provide opportunities for review, revision, and repetition, essential for memorizing formulas, principles, and problem-solving techniques, improving memory [18].
Numerous note-taking methods are in practice to make the learning process easy, effective, and time-saving. Unfortunately, there is no single method of note-taking that can serve the best for all. Different note-taking methods, their effectiveness, and taking the best notes have been the center of research for quite a long time. For example, see Refs. [31, 28, 25, 19, 32, 27, 20, 22, 21, 24, 23, 26, 29, 30].
A brief review of different note-taking methods is in order. Cornell note-taking is one of the most commonly used techniques of taking notes, in which each page is divided into four segments, namely, a line at the top for the topic, a few lines at the bottom for the summary, and two columns divided for the middle part of the paper with a narrow left for key points and wider right column for an extensive detail of the key corresponding points [33]. Significantly structured notes with main topic(s) followed by indented subtopic(s), sub-subtopic(s), and detailed elaborations underneath in a logical manner forming a skeleton of the textbook chapter, called outline method, serves as an excellent study guide when preparing for tests  [34]. One can also use the boxing method of note-taking, which involves drawing boxes around related pieces of information to visually organize and categorize thoughts, concepts, or ideas of the same category by visually separating different topics under different boxes and titles, making it easier to review and memorize the notes for longer [35]. The Zettelkasten method of note-taking is a personal knowledge management system, which serves well in creating an ever-growing web of knowledge over a certain topic, connecting fragmented note cards having one idea per note to a structured storing box, helping in the accumulation of knowledge on a specific topic over time [36]. In addition to being a memory retention tool, the Zettelkasten method fosters a deeper understanding of complex subjects and supports generating new ideas and insights, boosting creativity, productivity, and adaptability in organizing and processing information.
Figure 1: The figure shows an illustration of the mind mapping method about ways how students can increase attentiveness in the classroom to get a better learning experience.
One of the most sophisticated, dynamic, and visually stimulating note-taking methods, in which a learner writes the main idea in the center of the page and branches out with connecting lines to subtopics, supporting details, and related concepts, creating connections between them, is the mind mapping method [37]. Figure 1 illustrates a mind map of strategies that students can follow to increase attentiveness in the classroom. Some key strategies that students need to embrace to improve attentiveness are taking a digital detox, minimizing digital distractions, taking effective notes, actively participating in class, learning autonomy, properly managing tasks, and following a healthy lifestyle. The central concept of increasing attentiveness sits at the center and branches outwards to the six listed strategies, as shown in Fig. 1. The secondary, tertiary, and quaternary branches can be created, extending the key point to add sub-subtopics and details. Each strategy in Fig. 1 further branches outwards, connecting different approaches students may want to follow. For example, the note-taking strategy connects tactics of having handwritten notes instead of digital or online application written notes, adding visual concepts, using memory aids, preparing structurally consistent notes, and having mnemonics, symbols, and abbreviations instead of writing complete sentences or paragraphs.
A digital detox helps reduce stress by alleviating cognitive overload, improve memory, enhance focus, and allow the brain to rest and process information more effectively, thereby improving mental clarity [38]. Moreover, the digital detox helps in students’ psychological development and improvement of students’ concentration [39]. Although the use of digital technology is unavoidable in the current digital transformation era, a brief discussion on students’ necessity of taking a digital detox during classes to increase attentiveness is in order. Unless the learning strategy, instructional approach, and subject and content-like contextual factors demand, a wise choice for the students to improve attentiveness in class is not to use network and internet-connected electronic devices such as cell phones, smartwatches, laptops, tablets, etc., during lectures, which not only distract students but also negatively impact in students’ participation in classroom activities. Students are strongly recommended to turn off notifications during class and follow daily communication management strategies such as setting disconnection time, no screen time, communication preferred time, time for social media interaction, and decreasing internet addiction, etc. Mind mapping diagrams are free-flowing and helpful for brainstorming, finding linkage between ideas/concepts/topics, and generating ideas, keeping the main topic at the center. One may choose to prepare or recommend students to create concept maps as they are more structured, focused on connecting different concepts, and follow a hierarchy. Taking data from 55 studies about concept mapping in science achievement published in peer-reviewed journals and dissertations between 1980 and 2020, Anastasiou et al analyzed the effectiveness of concept maps science learning across several instructional settings, conditions, and methodological features [10]. Therein, the authors found that the concept maps were associated with increased science learning but there was a significant heterogeneity in most subsets of data.
Many other effective note-taking methods are also in practice such as the charting method, outline method, sentence method, flow method, boxing method, question-evidence-conclusion (QEC) method, split page method, writing on slides, bullet point method, list method, and rapid logging method [40, 41, 42, 43, 44, 45, 46]. The choice of note-taking method depends on many factors such as learning style, subject matter and topic, lecture structure and content delivery method, context, learner’s personal goal and preferences, available time to prepare notes, and need for mathematical, logical, and visual sophistication.
III Methods
III.1 Class and Lecture Design
The role of note-taking in learning introductory algebra-based undergraduate physics courses was studied in the minority-serving institutions, FSU and UNCP. The courses were designed based on the backward design of effective teaching, prioritizing the intended learning outcomes and course objectives over the chapters to be covered in the course, as well as the choice of instructional methods and assessment types.
A brief discussion on the backward design of instructional method is in order. The backward design instructional method is based on aligning instruction, course content, pedagogy, and assessment with the learning objective, in contrast to the traditional teaching method of starting the learning process based on teacher-designed content, delivering lectures, assigning problems, and assessing students based on what was taught [47]. In the backward design method, one first determines the desired learning objectives and goals that students are expected to learn by the end of the day, chapter, and the course. Second, assessments are determined and planned to check whether students have learned what they were expected to learn. Third, learning activities and tasks are prepared that will be done in the classroom and beyond to prepare students for assessments. Algebra-based physics is primarily designed for students majoring in life sciences rather than students majoring in physical science. Students pursuing careers in life sciences are often less motivated and less prepared for physics courses, which require active, goal-oriented learning rather than passive, teacher-centered methods. To address this, we used a backward design approach in teaching these courses.
The classes met twice a week for 75 minutes each day. Pedagogy includes lecturing, simulations, demonstrations, problem-solving, hands-on activities, think-pair-share, and class work as multiple choice questions. The classwork covered topics discussed at the beginning and end of the lecture, which helped ensure that students arrived on time and stayed for the whole class and boosted class attendance [13]. Students were strongly recommended to take a digital detox in the classroom, which minimizes digital distractions and helps minimize multi-tasking, which is considered one of the major learning pitfalls [48]. Students using electronic devices to take notes are prone to multitasking, but multitasking during learning impacts comprehension and knowledge retention as attention gets divided, focus decreases, the brain needs to pay the task-switching costs on refocusing, the brain receives cognitive load usually more than it can handle with no stress, and working memory is disrupted [49]. The assignments include online quizzes, mostly focused on concept development, and instructor-prepared homework assignments focused on problem-solving. The instructors used a liquid syllabus, which is a dynamic, web-based, interactive, and student-engaging syllabus that includes visual elements such as figures and videos, links to resources, and QR codes. The liquid syllabus has fixability in terms of dates, learning activities, and so on, to ensure the course objective is met by the end of the semester. The liquid syllabus, featuring more multimedia components, was adopted, inspired by the fact that students rarely read the syllabus once at the beginning of the semester and need to be made aware of the instructors’ expectations, requirements, and core objectives for the content and the course. Study shows liquid syllabus provides students with what they need to succeed in the course, increases equity, gives a sense of belonging and respect using student-friendly language and visuals, and provides students with a humanizing introduction to the course [50].
Students must have one prerequisite course to enroll in Physics-I at FSU and UNCP. At FSU, students must have a minimum passing grade in one math course out of 4 math courses college trigonometry, algebra and trigonometry, pre-calculus mathematics-I, or pre-calculus mathematics-II. At UNCP, students must have a minimum passing grade in one math course out of college algebra, pre-calculus, or calculus -I or II. Comparison of course contents and the academic standards show that the minimum criterion of the prerequisite for Physics-I is the same. In both institutions, successful completion of Physics-I is the prerequisite for Physics-II. To be noted, the general Physics course at FSU is equivalent to the college Physics course at UNCP.
III.2 Grading Policy
The grading practices (followed by Tables 1 and 2) were communicated at the beginning of the semester and frequently during the semester, and grade points are made student-trackable to ensure students understand the grading practices, can determine their weaknesses and strengths, and are focused on their goals. Assignments were structured to (i) reduce the negative impact of late work, keeping students responsible for submitting assignments on time, (ii) give chances to implement feedback for improvement, and (iii) provide opportunities for students to reflect on and improve their performance, which motivated student in learning by doing and revising. We focused on reducing the basis of grading by concentrating on academic achievement and excellence by excluding non-academic components such as attendance, seminar participation, group discussion, extra work for points, etc. Online tests were provided with three possible attempts for each question, in which a link to the relevant section in the eTextbook and available resources was provided. A hint with insight into how to answer the questions was made visible after the first attempt, and the detailed solutions were provided a minute after the deadline. Numerical values were randomized for each student, and the order of answer choices was shuffled in each attempt. The desired number of significant figures was made visible for each question. Each student got feedback on each attempt. The best score out of the maximum three attempts was recorded.
One of the bonafide meanings of learning is storing the received knowledge, concepts, and gained skills in the long-term memory, retrieving them effortlessly as needed, and applying them in relevant contexts. Senses first perceive knowledge, information, course contents, or any other learning materials, pass to the sensory memory, go to the working memory, and finally, it either gets lost or gets stored in the long-term memory based on the learner’s attention, elaborated rehearsal, and repetition [3, 4]. Students are encouraged to increase attentiveness, as we presented in Fig 1, and be active physically and mentally in the learning environment rather than just being passive listeners. Recognizing that elaborated rehearsal, implementation of feedback, repetition of tasks, and gaining overall skills are crucial, we assigned a lower weight to problem-solving homework assignments and online quizzes than terminal exams and the final comprehensive exam. For the same reason, the final exam had a slightly larger weight than each terminal exam.
FSU follows the FN grade policy, which means if students do not show up in the first week of the class, they will automatically drop out of the class after the instructor notifies the registration office. Students who dropped out from getting FN grades are not counted in the university census. UNCP also follows a similar policy. There was no mid-semester withdrawal in both courses in the FSU and UNCP, although some students did not complete many assignments, which impacted their grades. As students did not voluntarily withdraw and universities do not have a policy of dropping students because of such issues, these students’ scores are included in the analysis.
III.3 Study Design
Keeping the other factors the same, we implemented teaching pedagogy differing with the preparation and use of note-taking in two sets of sections. In each university, we taught effective note-taking methods, asked students to prepare handwritten notes, and let them use the notes during classwork, timed quizzes, and exams. Study shows that hand-written notes increase comprehension, create a personal connection to the learning materials, improve memory and recall of the facts, and activate brain memory areas more robustly than preparing notes using digital devices, showing the pen mightier than the Keyboard, although taking notes on electronic devices is faster, and the note taker can write to a larger extent of word-for-word what the teacher is saying in a class [51]. Students using electronic devices such as iPads, tablets, and laptops to take notes are usually prone to copying, pasting, and collecting information without a deep thought about what they are writing. In contrast, hand-written note-takers usually mentally process the information further before they write it down; while they write, they get their fingers’ extra muscle memory. Therefore, this research is designed as a comparative study of the effectiveness of handwritten notes in enhancing attention and effectiveness in physics learning. In order to maintain consistency, the quizzes and exams administered for both groups were problem-solving based, which demand critical thinking, creativity, and decision-making. Keeping all other conditions and class structure unaltered, the methodology was split into two main categories.
  1. 1.
    Handwritten notes are prohibited in exams; instead, a list of equations is provided:
      This policy was implemented in College Physics I (PHY 1500) at UNCP and General Physics I (PHY 115) at FSU, where the note-taking strategy was not discussed, nor were students allowed to use their notes in the quizzes and exams. However, a list of all required mathematical expressions covered in the chapters was provided during timed assignments and exams.
  2. 2.
    Handwritten note-taking is encouraged and permitted for use during exams:
      The note-taking policy was implemented in College Physics II (PHY 1510) at UNCP and General Physics II (PHY 116) at FSU. At the beginning of the semester, a few samples of the note-taking skeleton were provided, and instructors consistently observed the notes students were taking during the class for their effectiveness and guided them as per need. Only the handwritten notes prepared by students were permitted during quizzes and exams. The notes students used in the timed quizzes and exams were monitored.
In the section where the note-taking strategy was implemented, a few samples of skeleton notes of each kind discussed above, along with sample notes, were also provided in the very first week of the semester. In addition, students received timely feedback on their notes, and those who struggled with note-takings and other academic things were encouraged to use office hours for additional support. All instructors have at least eight regularly scheduled office hours per week to help students in various aspects, including course contents, their academic and career plans, and other academic support. Instructors asked students to make mind maps connecting different concepts in all the chapters covered as note-taking homework, which instructors found very helpful for students’ overall understanding of the course.
Table 1: FSU’s University catalog to assign the final letter grade for earned grade points on a scale of 0–100 and a GPA out of 4.0.
Points 90 – 100 80 – 89.9 70 – 79.9 60 – 69.9 $<$ 60
Grade A B C D F
GPA 4.0 3.0 2.0 1.0 0.0
Table 2: The UNCP’s University catalog to assign the final letter grade for earned grade points on a scale of 0–100 and a GPA out of 4.0.
Points 93–100 90–92.99 87-89.99 83-86.99 80-82.99 78-79.99 74-77.99 70-73.99 68-69.99 64-67.99 60-63.99 $<$ 60
Grade A A- B+ B B- C+ C C- D+ D D- F
GPA 4.0 3.7 3.3 3.0 2.7 2.3 2.0 1.7 1.3 1.0 0.7 0.0
IV Results and discussion
The grade that each student receives is the grade that students earn based on their academic performance, not measured against the performance of other students. The final letter grades are assigned according to the respective University’s catalogs, which are tabulated in Table 1 and 2. At FSU, using the traditional five-grade system, a letter grade of ”A” (GPA of 4.0 out of 4.0) is assigned if a student receives 90% of possible credits or more than that. Letter grades ”B” (GPA of 3.0 out of 4.0), ”C” (GPA of 2.0 out of 4.0), and ”D” (GPA of 1.0 out of 4.0) are assigned if a student’s grade point lies in the range of 80 to 89.9%, 70 to 79.9%, and 70 to 79.9%, respectively. A student fails the course if the student receives less than 60% of the total possible points. In the UNCP, the letter grade assigning catalog is not a traditional five-grade, but it follows a sophisticated letter grade and plus-minus system (see Table 2).
Figure 2: Bar chart showing letter grade percentage in sections with no note-taking implemented. If only the exam were considered, students’ grades and the final grade, incorporating all evaluating components included, would be presented in different colors. There were a total of 32 students: 20 from FSU and 12 from UNCP
Figure 3: Bar chart showing letter grade percentage in sections where students took notes and were allowed to use their notes during times assignments and exams. Students’ grades, if only the exam were considered and the final grade, incorporating all evaluating components included, would be presented in different colors. There were a total of 35 students: 23 from FSU and 12 from UNCP
Table 3: Statistical measures of student performance in two different approaches to teaching introductory physics. The highest, lowest, and average scores were presented on the scale of 0–100.
Note-taking policy ($\rightarrow$) $Withoutnote-taking$ $Withnote-taking$
[…]
Highest 96.596.596.5 96.7 97.597.597.5 99.1
Lowest 44.544.544.5 51.5 39.439.439.4 41.7
Average 70.070.070.0 79.9 76.676.676.6 83.7
Standard deviation 16.616.616.6 13.4 14.514.514.5 12.8
At UNCP, each letter grade comes with three variations of a letter, followed by plus, letter only, and letter followed by a minus sign, except for the letter grade ”A,” which has only two variations of A and A-, and letter grade F (the failing grade) with no variation. Note that the plus and minus designation window for letter grades B and C are not the same, while the distribution range is similar for C and D. As the collective grade point range associated with a letter grade and its plus and minus at UNCP is the same to that of the letter grade at FSU, for a consistent analysis, we counted each letter along with its plus and minus at UNCP as the corresponding letter grade’s frequency such as sum of A- and A’s frequencies at UNCP is equivalent to A’s frequency at FSU, sum of B-, B+, and B’s frequencies at UNCP is equivalent to B’s frequency at FSU, and so on. Figures 2 and 3 show the percentage of letter grades that students received in sections with no note-taking and sections with note-taking were implemented. he sample size for these two groups in the study is almost the same, one being 32 and the other being 35. To better understand the effect of effective note-taking, the grades students receive when only exam scores are considered in assigning grades are presented separately. This analysis shows that homework, online quizzes, class quizzes, and other assignments helped improve students’ grades. Almost one-third of the students would fail in the sections with no note-taking implemented if only exams were considered while assigning the letter grade. Yet, nearly one-tenth of the students failed the course, even with the inclusion of all grading components. The data show that when students were allowed to take notes, the percentage of students receiving better letter grade (by better letter grade, we mean ”A”) did not change significantly. The percentage of students getting the letter grade ”A” is almost the same (please see and compare the pair of bars for letter grade ”A” of Fig. 2 with the same of Fig. 3) . The slight percentage difference is primarily due to the fact that the sample sizes are close, but not identical. While the same is not true for the percentage of students getting the letter grade ”B” and lower. With note-taking, exam-only grade follows the bell curve, peaking at B and having a tail towards a lower grade. The letter grade significantly improved when students were allowed to prepare and use the notes they had prepared. Furthermore, the percentage of students who failed the course reduced to 2.9% only. Our observation is that note-taking helped average-performing (students getting B and C letter grades), less motivated, and less interactive students more than it did comparatively better-performing students.
[…]
The universities involved in the study are constituent universities of the same university system with similar standards. The instructors of the two institutions coordinated on content taught, learning activities, assessments, and grading policies and standards. Students’ scores in each assignment were collected in Canvas and shared among instructors. The grade sheets can be exported and made available, with students’ personal information redacted. Statistical calculations were performed using Excel and Python packages. The findings of this study, concluding that effective note-taking helps in academic excellence, are consistent with those of other studies examining the role of effective note-taking in learning.
V Statistical Tests and Limitation of the Study
We performed a $\chi^{2}$ goodness-of-fit test to determine whether the distributions of exam-only scores and aggregated scores were statistically similar. The scores were grouped into custom ranges: below 60, 60–70, 70–80, 80–90, and 90–100, and the $\chi^{2}$ values were evaluated using these categories as
$\chi^{2}=\sum_{i=1}^{n}\frac{(E_{i}-A_{i})^{2}}{E_{i}}\,,$ (1)
where $E_{i}$ and $A_{i}$ are frequencies for exam only and aggregated scores in $i^{\rm th}$ interval. The $\chi^{2}$ values for Phys-I and Phys-II are 23.9 and 24.5 respectively, indicating that the frequency distributions differ significantly. The $p$-values in each case is less than 0.05 suggesting that the difference in these two distributions in each course is statistically significant and they did not occur due to random chance. At a significance level of $\alpha=0.05$ (or 95% confidence interval), the statistical power is about 70%, meaning there is a 70% chance of correctly detecting a true effect. In statistical hypothesis testing, this is known as a Type II error, also called a false negative [52].
Figure 4: Plot showing the required sample size for 80% statistical power for as a function of effect size in terms of Cramér’s $V$-values. For the $V$-values of the students’ grade sheets presented in this study, the required sample size for Phys-I to get 80% statistical power is 39 and that for Phys-II is 45, which are slightly larger than the sizes in this study.
To gain a better understanding of the results, one can evaluate the effect size using measures like Cramér’s V, which can be calculated using the following formula:
$V=\sqrt{\frac{\chi^{2}}{n\cdot(k-1)}}\,,$ (2)
where $n$ and $k$ are total number of observations and number of categories respectively. The Cramér’s $V$ values in Phys-I data is 0.45 while that for Phys-II is 0.42. To reliably reject a false null hypothesis, a statistical power of 80% or higher is commonly recommended at a significance level of $\alpha=0.05$ [53]. To increase reliability, a larger sample size would help.
The study was conducted at only two institutionseach offering a limited number of introductory physics course sections. Expanding the study to include multiple institutions and a larger student population would enhance the reliability of the results. Reliability could be improved by conducting the study with half of the students in a class using note-taking strategies, while the other half do not, in which students could self-select their group or be randomly assigned by the instructor. Instead, one can implement these two strategies in different chapters/modules for a chosen course/section. Alternatively, one can investigate the role of note-taking by implementing the above-mentioned strategies to assess the same group of students across different midterm exams.
VI Conclusion
The impact of effective note-taking on learning in undergraduate-level algebra-based introductory physics courses at minority-serving universities was studied. We revisited the importance of a distraction-free and student-centered active learning environment to enhance the retention of learning materials in long-term memory, which students can retrieve effortlessly as needed and apply the received knowledge, concepts, and gained skills in various contexts now and in the future. A few commonly used note-taking methods were discussed. As an example of the mind mapping method of note-taking, we have presented a mind map to discuss techniques that students may follow to increase their attentiveness in the classroom, making physics learning fruitful. We discussed various approaches we have implemented in teaching and assessing algebra-based undergraduate-level introductory physics courses, as well as the impact of our efforts to make physics learning fun, active, student-centered, interactive, and objective-oriented.
We found that the majority of students were motivated to prepare better notes and actively engage in the classroom, as they could use their handwritten notes in timed assignments and exams. The final letter grades that students received at the end of the semester were recorded as a quantitative measurement of learning effectiveness. The letter grade significantly improved when students were allowed to prepare and use the notes they prepared during exams and timed assignments. Not only did the percentage of students getting higher grades increase with note-taking implementation, but the percentage of students failing the course also decreased sharply. The study revealed that implementing a note-taking learning approach comparatively enhances the grades of low—and average-performing students more than those of high-performing students.
The findings recommend that educators consider effective note-taking as an integral part of learning activities. In the current age of digital distraction, it is more important than ever. Asking students to make their mind map connecting different concepts in all the modules (chapters) covered in a course as note-taking homework, which we encouraged students to prepare, is excellent homework any instructor can assign. The mind mapping diagram presented in the manuscript serves as a strong example.